# Decouple and Decompose: Scaling Resource Allocation with DeDe

### v2

@@ 8 Related Work
Resource allocation problems.  Multi-tenant systems often address resource allocation challenges. Section 5 examines three prominent examples: cluster scheduling [46, 27, 44, 61, 64] allocates jobs to CPUs or GPUs in heterogeneous clusters, traffic engineering [24, 1, 67, 50, 31] allocates traffic demands to network link capacities in wide-area networks, and load balancing [16, 48, 52] allocates query loads to servers in distributed databases. Beyond these problems, Skyplane [26] allocates cloud resources (e.g., bandwidth and TCP connections) to overlay paths that traverse intermediate cloud regions, minimizing the cost of inter-region bulk transfers. Zeta [65] allocates traffic forwarding capability from gateway clusters to virtual private clouds for scalable and robust east-west communication in large-scale clouds. Shoofly [51] allocates light wavelengths of optical fiber to network shortcuts abstracted from optical bypasses, in order to minimize the hardware costs of provisioning long-haul capacity.
@@
Optimization algorithms.  Various constrained optimization algorithms have been employed to solve resource allocation problems, depending on the variable types, objectives, and constraints. For linear programming problems like traffic engineering, the Gurobi solver [22] utilizes two primary methods. The simplex method [45] iteratively progresses along the boundaries of the feasible region toward the optimal solution, while the barrier method [30] iteratively approaches the optimal solution from within the feasible region. In mixed-integer linear programming like load balancing, the branch-and-bound method [8] divides the problem into smaller subproblems, or branches, and then eliminates certain branches based on bounds on the optimal solution at each iteration. For more complex optimization problems, such as cone programming when maximizing proportional fairness in cluster scheduling, the ECOS algorithm [18] adopts a similar iterative process as the barrier methods, with search directions found by a symmetric indefinite KKT system. All these constrained optimization algorithms operate sequentially and iteratively, and thus introduce bottlenecks in solving time.
@@
Decomposition for scalability.  As resource allocation problems grow, efforts have been made to parallelize the iterative process in optimization algorithms for better scalability. NCFlow [1] accelerates traffic engineering by partitioning the network into clusters and concurrently solving the flow allocation problem in each cluster. POP [43] targets “granular” allocation problems where resources are interchangeable. It randomly splits resources and demands into subsets and allocates each subset in parallel. Soroush [42] generalizes the classical waterfilling algorithm for max-min fair allocation and presents parallelizable combinatorial algorithms with fairness guarantees. Teal [63] models traffic engineering with neural networks to benefit from the massive parallelism in GPU. Eason et al. [19] employ the Benders decomposition algorithm to parallelize the design of a specialized cross-layer backbone network, offering performance guarantees for a particular allocation problem.
@@ 9 Conclusion
In this paper, we scale real-world resource allocation problems from a fresh perspective, observing that most of them exhibit an inherent separable structure. This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands. We implement this approach, DeDe, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation. Through extensive evaluation across diverse resource allocation tasks, we demonstrate that DeDe significantly improves scalability while closely approximating the quality of exact solutions.
Appendix A Mathematical Representation of DeDe
A resource allocation problem can be written as
$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(x_{*j}),$ (1)
$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
$D_{j}x_{*j}=d_{j},\quad\forall j,$ (3)
where Equation 1 is the objective, and Equation 2 and Equation 3 are resource constraints and demand constraints, respectively.
Our first step decouples resource and demand constraints. We create a copy of the allocation variable $x$, denoted as $z$, and replace the variable $x$ in demand constraints (Equation 3) as well as the demand-related portion of the objective function (Equation 1) with the variable $z$. We also add a new constraint $x=z$ (Equation 6) to keep $x$ and $z$ equal.
$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}),$ (1)
$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
$D_{j}z_{*j}=d_{j},\quad\forall j,$ (3)
$x-z=0.$ (6)
We rewrite our problem in the Augmented Lagrangian to move the constraints to the objective function:
$\begin{split}&\mathcal{L}_{\rho}(x,z,\alpha,\beta,\lambda)=\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j})\\ &+\frac{\rho}{2}\Bigl{(}\sum_{i}||R_{i}x_{i*}-r_{i}+\alpha_{i}||_{2}^{2}+\sum_{j}||D_{j}z_{*j}-d_{j}+\beta_{j}||_{2}^{2}\\ &+||x-z+\lambda||_{F}^{2}\Bigr{)}-\frac{\rho}{2}(||\alpha||_{2}^{2}+||\beta||_{2}^{2}+||\lambda||_{F}^{2}).\end{split}$
ADMM allows partial optimization alternatively on $x$ and $z$:
$\alpha_{i}^{(k+1)}$ $:=\alpha_{i}^{(k)}+R_{i}x^{(k+1)}_{i*}-r_{i},\forall i$
$\beta_{j}^{(k+1)}$ $:=\beta_{j}^{(k)}+D_{j}z^{(k+1)}_{*j}-d_{j},\forall j$
$\lambda^{(k+1)}$ $:=\lambda^{(k)}+x^{(k+1)}-z^{(k+1)}$
Our second step further decomposes the partial optimization into optimizing each row in $x$ and each column in $z$. We rewrite Equation 7 as
$+\frac{\rho}{2}\big{\|}x-z^{(k)}+\lambda^{(k)}\big{\|}_{F}^{2}.$
We rearrange the parentheses in Equation 2 to make the variables in each parenthesis mutually exclusive:
$x^{(k+1)}=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{i}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
With the variable in each parenthesis mutually exclusive, we decompose to optimize per-resource allocation $x_{i*}$ individually:
$\begin{split}x_{i*}^{(k+1)}=&\operatorname*{arg\,min}_{x_{i*}\in\mathcal{X}_{i*}}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.\end{split}$
We follow a similar decomposing procedure for demand-constrained optimization on $z$:
$z^{(k+1)}=$ $\operatorname*{arg\,min}_{z}\mathcal{L}_{\rho}(x^{(k+1)},z,\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (8)
$=$ $\operatorname*{arg\,min}_{z}\sum_{j}g_{j}(z_{*j})+\frac{\rho}{2}\sum_{j}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}$
$+\frac{\rho}{2}\big{\|}x^{(k+1)}-z+\lambda^{(k)}\big{\|}_{F}^{2}$
$=$ $\operatorname*{arg\,min}_{z}\sum_{j}\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}$
$+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.$
The optimization of $z$ can now be decomposed into optimization on per-demand allocation $z_{*j}$:
$\begin{split}z_{*j}^{(k+1)}=&\operatorname*{arg\,min}_{z_{*j}}\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.\end{split}$

### v3

@@ 8 Related Work
Resource allocation problems.  Multi-tenant systems often address resource allocation challenges. Section 5 examines three prominent examples: cluster scheduling [47, 28, 45, 63, 66] allocates jobs to CPUs or GPUs in heterogeneous clusters, traffic engineering [24, 1, 69, 51, 32] allocates traffic demands to network link capacities in wide-area networks, and load balancing [14, 49, 53] allocates query loads to servers in distributed databases. Beyond these problems, Skyplane [27] allocates cloud resources (e.g., bandwidth and TCP connections) to overlay paths that traverse intermediate cloud regions, minimizing the cost of inter-region bulk transfers. Zeta [67] allocates traffic forwarding capability from gateway clusters to virtual private clouds for scalable and robust east-west communication in large-scale clouds. Shoofly [52] allocates light wavelengths of optical fiber to network shortcuts abstracted from optical bypasses, in order to minimize the hardware costs of provisioning long-haul capacity. By leveraging the commonly observed separable structure of real-world resource allocation problems, DeDe is applicable to the aforementioned cases as well as a broad range of others.
@@
Optimization algorithms.  Various constrained optimization algorithms have been employed to solve resource allocation problems, depending on the variable types, objectives, and constraints. For linear programming problems like traffic engineering, the Gurobi solver [22] utilizes two primary methods. The simplex method [46] iteratively progresses along the boundaries of the feasible region toward the optimal solution, while the barrier method [31] iteratively approaches the optimal solution from within the feasible region. In mixed-integer linear programming like load balancing, the branch-and-bound method [7] divides the problem into smaller subproblems, or branches, and then eliminates certain branches based on bounds on the optimal solution at each iteration. For more complex optimization problems, such as cone programming when maximizing proportional fairness in cluster scheduling, the ECOS algorithm [16] adopts a similar iterative process as the barrier methods, with search directions found by a symmetric indefinite KKT (Karush-Kuhn-Tuck) system. All these constrained optimization algorithms operate sequentially and iteratively, and thus introduce bottlenecks in solving time. In contrast, DeDe decouples and decomposes the problem into many parallel subproblems, and thus allows faster problem solving.
@@
Decomposition for scalability.  As resource allocation problems grow, efforts have been made to parallelize the iterative process in optimization algorithms for better scalability. NCFlow [1] accelerates traffic engineering by partitioning the network into clusters and concurrently solving the flow allocation problem in each cluster. POP [44] targets “granular” allocation problems where resources are interchangeable. It randomly splits resources and demands into subsets and allocates each subset in parallel. Soroush [43] generalizes the classical waterfilling algorithm for max-min fair allocation and presents parallelizable combinatorial algorithms with fairness guarantees. Teal [65] models traffic engineering with neural networks to benefit from the massive parallelism in GPU. Eason et al. [17] employ the Benders decomposition algorithm to parallelize the design of a specialized cross-layer backbone network, offering performance guarantees for a particular allocation problem. While these prior decompositions rely on restrictive assumptions or problem-specific designs, DeDe offers a general and theoretically grounded approach solution for diverse resource allocation problems.
@@ 9 Conclusion
In this paper, we scale real-world resource allocation problems from a different lens, observing that the vast majority of them exhibit an inherent separable structure. This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands. We implement this approach, DeDe, as a Python package, offering a push-button solution that accelerates large-scale resource allocation. Through extensive evaluation across diverse resource allocation tasks, we demonstrate that DeDe significantly improves scalability while closely approximating the quality of exact solutions.
Acknowledgments
We thank our anonymous reviewers and shepherd for their insightful comments. We also thank Yang Zhou, Justin Chiu, Alexander Rush, and Victor Bahl for their valuable feedback and support. We extend special thanks to Suvan Chatakondu for enhancing our code after the paper’s acceptance. This work is supported in part by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. This work is also supported by National Science Foundation grant CCF-2326605 (Expedition in Computing).
Appendix
Appendix A Evaluation Setup of Cluster Scheduling
We detail the experimental setup for the large-scale cluster scheduling evaluation described in §7.1.1.
Computing resources.  We collect 456 distinct GPU/CPU resource types from three benchmark sources [26, 36, 19]. These resource types differ along multiple dimensions such as vendor, interconnect, memory configuration, deployment platform, etc. While it is unlikely today for any single cluster to host all 456 types, recent trends in distributed ML workloads [66, 62] suggest the possibility of scheduling ML jobs across global-scale clusters in the future. For each resource type, the number of available instances is randomly drawn from the set $\{8,16,24,\ldots,64\}$, reflecting common modern hardware configurations where nodes are provisioned with GPU/CPUs in multiples of eight. This setup results in a total of 16,520 GPU/CPU instances in our simulated environment.
Jobs.  We generate 2,588 types of ML jobs based on the Notable AI Models dataset [18]. For each ML model, we synthesize multiple job types encompassing both training and inference tasks under various numerical precisions. The number of compute instances requested by each job is sampled from $\{1,2,4,8,16,32\}$, in line with common practices in data and model parallelism. Job throughput values are either taken directly from ML benchmarks [26, 36, 19] or, when unavailable, estimated based on each job’s FLOP requirements and the computational capacity of the respective hardware.
Scheduling Simulator.  Our evaluation employs Gavel’s ML scheduling simulator [45], with several adjustments. Jobs are time-sliced across available resource types but are not co-located, due to the increased size and resource demands of modern ML workloads. To ensure that jobs do not grow beyond cluster capacity, we set the average job inter-arrival time to 100 seconds in the Poisson process. We begin the simulation from a steady state, maintaining approximately 1,500 to 2,000 active jobs at any given time. Job allocation decisions are made every 6 minutes, consistent with Gavel’s default settings, and the simulation runs for 200 scheduling rounds, or 20 hours.